OpenStudy (anonymous):

A can of Pepsi is supposed to contain, on the average, 12 ounces of soda with a standard deviation of 0.3 ounces. Suspecting fraud, you take a random sample of 40 cans and measure the amount of Pepsi in each. Your measurements show that the 40 cans had a mean of11.9 ounces. What is the probability of a random sample of 40 cans having a mean of 11.9 ounces of soda or less?

OpenStudy (anonymous):

@jim_thompson5910 this is my last question.

jimthompson5910 (jim_thompson5910):

are you familiar with the xbar distribution?

OpenStudy (anonymous):

It sounds familiar but I don't remember. How do I do that?

jimthompson5910 (jim_thompson5910):

have you learned about the normal distribution at all?

OpenStudy (anonymous):

Uhh, not that I really remember. I don't know.

jimthompson5910 (jim_thompson5910):

ok one sec

ganeshie8 (ganeshie8):

heard of `Central Limit Theorem` ?

OpenStudy (anonymous):

Nope. What's that?

ganeshie8 (ganeshie8):

you need to use CLT to get the normal distribution parameters, and find the p-value for 11.9

ganeshie8 (ganeshie8):

First list down the given info : \(\mu = 12\) \(\sigma = 0.3\) \(n=40\)

ganeshie8 (ganeshie8):

start by finding the standard error : \(SE = \dfrac{\sigma}{\sqrt{n}} = ?\)

OpenStudy (anonymous):

0.3/(srt40) = 0.04743416491 is that right?

ganeshie8 (ganeshie8):

thats right ! 0.3/sqrt(40) is ur new standard deviation

ganeshie8 (ganeshie8):

CLT tells u that the new mean is same as population mean = 12

ganeshie8 (ganeshie8):

next find the zscore for 11.9

ganeshie8 (ganeshie8):

heard of `zscore` before right ?

OpenStudy (anonymous):

How do I get zscore?

ganeshie8 (ganeshie8):

\(Zscore = \dfrac{\text{observation - mean}}{\text{standard deviation}} = \dfrac{11.9 - 12}{\frac{0.3}{\sqrt{40}}}\)

ganeshie8 (ganeshie8):

Once you have the Zscore, lookup in Ztable for the corresponding probability

ganeshie8 (ganeshie8):

http://www.math.upenn.edu/~chhays/zscoretable.pdf

OpenStudy (anonymous):

I got -2.108185106

ganeshie8 (ganeshie8):

thats right ! -2.108185106 is ur Zscore

ganeshie8 (ganeshie8):

to get the probability for that Zscore, u need to lookup in Ztable

ganeshie8 (ganeshie8):

http://www.math.upenn.edu/~chhays/zscoretable.pdf

ganeshie8 (ganeshie8):

know how to interpret Ztable ?

OpenStudy (anonymous):

I don't understand how the ztable works...

ganeshie8 (ganeshie8):

its easy, u familiar wid microsoft excel ?

OpenStudy (anonymous):

a little

ganeshie8 (ganeshie8):

a little is enough :)

OpenStudy (anonymous):

alright :) so how does it relate to the ztable?

ganeshie8 (ganeshie8):

-2.108185106 round it to two decimal places : -2.11

ganeshie8 (ganeshie8):

open this table : http://www.math.upenn.edu/~chhays/zscoretable.pdf

OpenStudy (anonymous):

0.0174?

ganeshie8 (ganeshie8):

You got it !

OpenStudy (anonymous):

Really? Yay!

ganeshie8 (ganeshie8):

just fix ur one finger at ROW starting wid `-2.1` and another finger at column having `0.01` the intersection is ur probability : `0.0174`

ganeshie8 (ganeshie8):

yup ! thats all to Ztables :)

ganeshie8 (ganeshie8):

Ztable gives the probability (area under normal curve) for <= Zscore (gives the area for left tail)

ganeshie8 (ganeshie8):

if above statement confuses, dont bother about it..

ganeshie8 (ganeshie8):

0.0174 is the `probability for having a sample of 40 cans whose mean is 11.9`

ganeshie8 (ganeshie8):

0.0174 is same as 1.74 % which is pretty less...

ganeshie8 (ganeshie8):

so, looks like there is indeed some fraud going on...

OpenStudy (anonymous):

Really? So that's the answer to the question?

ganeshie8 (ganeshie8):

whats the question ?

ganeshie8 (ganeshie8):

this is ur question : What is the probability of a random sample of 40 cans having a mean of 11.9 ounces of soda or less?

ganeshie8 (ganeshie8):

0.0174 is the probability you're done.

ganeshie8 (ganeshie8):

if u prefer percentages, u can say the probability is 1.74 % either way is fine

OpenStudy (anonymous):

Really? Thank you! I really appreciate your help :)

ganeshie8 (ganeshie8):

np :) usually these questions are closely related to hypothesis testing... where they ask u figure out whether the fraud is going on or not...

ganeshie8 (ganeshie8):

for hypothesis testing - u need to work it the same way, after finding the probability u need to make a decision also.

ganeshie8 (ganeshie8):

for example, here since the probability is very less, u might make a decision in favor of the claim that the fraud exists.